x^2+y^2-9=0
circle ellipse parabola or hyperbola
???

Answer by user_dude2008(1862)

x^2+y^2-9=0 ----> x^2+y^2=9

eqn fits form: x^2+y^2=r^2 (circle)


Ans: x^2+y^2-9=0 is a circle